OpenStudy (anonymous):

help! What would the solution set of this system of inequalities look like for this problem? (Will post in a moment)

OpenStudy (anonymous):

\[y \ge \frac{ -x }{ 2 } + 2\] \[y \le x+ 4\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@pooja195